OpenStudy (anonymous):

5v√ 75vw^2 + w√ 27v^3 simplify

OpenStudy (anonymous):

anybody know?

OpenStudy (campbell_st):

simplify both radicals... \[5v \sqrt{25w^2 \times 3v} + w \sqrt{9v^2 \times 3v}\] this can be written as \[5v \sqrt{25w^2} \times \sqrt{3v} + \sqrt{9v^2} \times \sqrt{3v}\] hopefully this helps... you can take the square root of the perfect squares... and you'll be teft with the same radical that can be added.

OpenStudy (anonymous):

im lost as heck bro

OpenStudy (campbell_st):

ok... so I split 75vw^2 into 2 perfect squares 25 x 3 = 75 v x w^2 = vw^2 so 75 vw^2 can be written as 25w^2 x 3v does that make sense...?

OpenStudy (anonymous):

ya

OpenStudy (campbell_st):

ok... so you are happy to take the square root of 25w^2..? the same thing has been donw with the 2nd radical 27 = 9 x 3 v^3 = v^2 x v so 27v^3 = 9v^2 x 3v take the square root of 9v^2..... the 3v in both cases will remain in the radical.